OpenStudy (anonymous):

dibbie anderson a piece of wood measuring 3 5/8 inches into two equal pieces. howlong is each piece? the length of each piece is

OpenStudy (campbell_st):

convert it to an improper fraction so you get \[3\frac{5}{8} = \frac{3 \times 8 + 5}{8} = \frac{29}{8}\]\ so dividing by 2 is the same as multiplying by a half. so you have \[\frac{29}{8} \times \frac{1}{2} = \frac{29 \times 1}{8 \times 2}\] just simplify and change back to a mixed number

OpenStudy (whpalmer4):

Alternatively, divide 3 by 2 = 3/2 divide 5/8 by 2 = 5/16 add 3/2 + 5/16 =

OpenStudy (anonymous):

the piece of wood is called dibbie anderson?

OpenStudy (anonymous):

im lost

OpenStudy (anonymous):

lol sorry

OpenStudy (anonymous):

ok 3 5/8 and 3 5/8

OpenStudy (whpalmer4):

No, 1 piece of wood, 3 5/8 inches long. Cut in two equal lengths, how long is each length? You can convert 3 5/8 to an improper fraction (\(\frac{3*8+5}{8} = \frac{29}{8}\)) and divide the fraction by 2 (\(\frac{29}{16}\)) and convert back to a mixed fraction (\(1 \frac{13}{16}\)) Or, you can divide each component of the mixed fraction in half: 3 divided by 2 (3/2) 5/8 divided by 2 (5/16) add together (\(\frac{3}{2} + \frac{5}{16} = \frac{24}{16} + \frac{5}{16} = \frac{29}{16}\)) convert back to mixed fraction as above (\(1 \frac{13}{16}\)) This works because \[\frac{a + b}{2} = \frac{a}{2} + \frac{b}{2}\]

OpenStudy (whpalmer4):

Which is easier depends on your proficiency with fractions, and the particular fractions in the problem.